英語 (English)

発信型英語 (Communicative English) (医 ((看 a)2 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 2 単位後期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This Communicative English class considered to provide very good

aid in growing up English communicating abilities. 【授業の概要】Classes will be running in ENGLISH language only for more

practice chance. There is no specific text book but there will be some other media depending on the ’Topic’ which will be selected up to date. Listening and Conversational practices will take place during the classes.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 To build up more self-confidence and higher effective communication

skills. 【授業の計画】

1. Overall Guidance

2. Selected Conversation 1

3. Selected Conversation 2

4. Selected Conversation 3

5. Selected Conversation 4

6. Selected Conversation 5

7. Selected Conversation 6

8. Selected Conversation 7

9. Selected Conversation 8

10. Selected Conversation 9

11. Selected Conversation 10

12. Selected Conversation 11

13. Selected Conversation 12

14. Selected Conversation 13

15. Selected Conversation 14

16. Review 【教科書】There is no specific text book but there will be some other media

depending on the ’Topic’ which will be selected up to date.

【参考書等】 [参考資料]

【成績評価の方法】 Attendance, Homework fulfillment and Participation in class. 【再試験の有無】 There is no make-up examination for this course. 【受講者へのメッセージ】 Asking questions in class is very important.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221211 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 座喜 .

English

Communicative English (医 ((看 a)2 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 2 units 後期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This Communicative English class considered to provide very good aid in

growing up English communicating abilities. Outline⟩ Classes will be running in ENGLISH language only for more practice

chance. There is no specific text book but there will be some other media depending on the ’Topic’ which will be selected up to date. Listening and Conversational practices will take place during the classes.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ To build up more self-confidence and higher effective communication skills.

Schedule⟩

1. Overall Guidance

2. Selected Conversation 1

3. Selected Conversation 2

4. Selected Conversation 3

5. Selected Conversation 4

6. Selected Conversation 5

7. Selected Conversation 6

8. Selected Conversation 7

9. Selected Conversation 8

10. Selected Conversation 9

11. Selected Conversation 10

12. Selected Conversation 11

13. Selected Conversation 12

14. Selected Conversation 13

15. Selected Conversation 14

16. Review Textbook⟩ There is no specific text book but there will be some other media

depending on the ’Topic’ which will be selected up to date. Reference⟩ [参考資料] Evaluation Criteria⟩ Attendance, Homework fulfillment and Participation in class. Re-evaluation⟩ There is no make-up examination for this course. Message⟩ Asking questions in class is very important.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221211 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki .